# Multi-Continental Healthcare Modelling Using Blockchain-Enabled Federated Learning

### v1

@@ I Introduction
In this paper we propose Multi-Continental Glucose Prediction (MCGP) framework using blockchain-enabled federated learning. This technique allows participating organizations to contribute to model training without the need to share sensitive health data directly. Instead, model parameters are exchanged among participants, thus preserving the privacy of the underlying health data. It also has protocols to apply incentive to encourage good contributors and to build robustness in detecting malicious behaviors. For the first time, We tested proposed modeling framework within a network spanning three continents—Europe, North America, and Asia—demonstrating that the model derived from this collaborative approach outperforms those trained on isolated datasets from individual health data, shown in Fig 1. This framework not only proves effective for blood glucose prediction but can also be readily extended to model other chronic diseases or address broader healthcare challenges.
@@ III-B Federated Learning for Glucose Prediction
Initially, all hospitals and the central server initialize a training model, denoted as $\theta^{r-1}$. The local models at each hospital are aligned with the global model on the central server to ensure a consistent starting point. In the first training round ($r=1$), each hospital uses its own training data to update its local model $\theta^{r-1}_{k}$. After a predefined number of epochs ($E$), the updated local model weights and the number of data samples ($n_{k}$) are sent to the central server. The central server aggregates the updated model weights from all hospitals using a weighted averaging approach:
@@ III-C1 Blockchain and Smart Contract
Blockchain is a decentralized and secure ledger technology that ensures data integrity and transparency across multiple nodes. Smart contracts are self-executing agreements coded directly onto the blockchain, automating processes and enforcing terms without intermediaries. In BCFL, blockchain and smart contracts can play a crucial role by providing a secure and transparent framework for decentralized machine learning. Blockchain can ensure the integrity and traceability of data and model updates, while smart contracts can automate the coordination and validation of contributions from different parties. For instance, smart contracts can be used to implement incentive mechanisms for participants in the blockchain-based FL.
@@ III-C2 Blockchain Tokens
In our BCFL system, a blockchain token, named as \$FML, serves as a quantitative measurement of assets, providing a reliable indicator of a participant’s trustability within our system. By staking \$FML tokens, participants demonstrate their commitment and reliability, as their participation is backed by a tangible, verifiable asset. Their tokens will be integrated into our incentive mechanism, which not only incentivizes honest behavior but also deters malicious activities, as participants risk losing their staked tokens if they act dishonestly. Similar to ERC20 tokens, the \$FML tokens in our BCFL system are fungible and governed by smart contracts. We assume that all participants have acquired sufficient \$FML tokens (e.g., by purchasing \$FML from centralized or decentralized exchanges) before joining the BCFL system.
@@ III-C3 On-Chain Incentive Mechanism
As shown in Figure 3, we follow the existing work [17] and adopt an incentive mechanism to reward honest behaviors and penalize malicious behaviors.
1. Participants need to stake a fixed amount (i.e., $\delta$) of \$FML tokens into the smart contract to be eligible to participate in the FL training process. 2. In each round, a participant will be randomly selected to act as either an FL training node or a voter by using a verifiable random function (VRF) [19]. 3. A training node is responsible for performing local training using its own data and uploading the local model updates to a public database (e.g., IPFS [^1] https://ipfs.tech/). 4. Voters aggregate all the local model updates for that round to obtain the global model update. They evaluate the global model by running it on their evaluation data to propose voting results, i.e., $\mathsf{vote}_{i}^{r}\in\{\texttt{support},\texttt{oppose}\}$. 5. Based on the majority of the voting result (i.e., $\mathsf{majorVote}^{r}\in\{\texttt{support},\texttt{oppose}\}$), participants in that round will either be rewarded or have their FML tokens slashed. Specifically, if a voter’s choice, $\mathsf{vote}_{i}^{r}$ differs from the majority result (i.e., $\mathsf{vote}_{i}^{t}\not=\mathsf{majorVote}^{r}$), that voter will be penalized; otherwise, they will be rewarded. For all selected training nodes in this round, the majority voting result will determine whether they receive a reward or not. 6. At the beginning of each round, we check the participants’ remaining \$FML token balances in the smart contract: if their balance is still larger than a predefined threshold (i.e., $\delta_{0}$), they will be eligible to continue the following round. Otherwise, they will be removed from the system. 7. For each round $r$, the finalized aggregated global model update is determined by the majoriting voting result: $\theta^{t}=\begin{cases}\hat{\theta^{r}},&\mathsf{majorVote}^{r}=\texttt{support}\\ \theta^{r-1},&\mathsf{majorVote}^{r}=\texttt{oppose}\end{cases}$
@@
As shown in [17], the expected return for malicious participants is negative, ultimately resulting in their expulsion from the system. Different from the existing work [17], which utilizes smart contracts to perform on-chain aggregation of local model updates, our approach allows voters to execute the aggregation off-chain. This improvement can significantly reduce on-chain costs (e.g., gas fees), particularly when dealing with training models that have a large number of parameters.

### v2

@@ I Introduction
In this paper, we propose a Multi-Continental Glucose Prediction (MCGP) framework using blockchain-enabled federated learning. This technique allows participating organizations to contribute to model training without the need to share sensitive health data directly. Instead, model parameters are exchanged among participants, thus preserving the privacy of the underlying health data. It also has protocols to apply incentive mechanisms to encourage good contributors and to build robustness in detecting malicious behaviors. For the first time, we tested the proposed modeling framework within a network spanning three continents—Europe, North America, and Asia—demonstrating that the model derived from this collaborative approach outperforms those trained on isolated datasets from individual health data, shown in Fig 1. This framework not only proves effective for blood glucose prediction but can also be readily extended to model other chronic diseases or address broader healthcare challenges.
@@ III-B Federated Learning for Glucose Prediction
Initially, all hospitals and the central server initialize a training model, denoted as $\theta^{\prime}$. The local models at each hospital are aligned with the global model on the central server to ensure a consistent starting point. In the first training round ($r=1$), each hospital uses its own training data to update its local model $\theta_{k}^{\prime}$. In subsequent rounds ($r>1$), each hospital updates its local model starting from the previous round’s aggregated global model $\hat{\theta}^{r-1}$. After a predefined number of epochs ($E$), the updated local model weights $\theta_{k}^{r}$ and the number of data samples ($n_{k}$) are sent to the central server. The central server aggregates the updated model weights from all hospitals using a weighted averaging approach:
@@ III-C1 Blockchain and Smart Contract
Blockchain is a decentralized ledger technology that ensures data integrity and transparency across multiple participants in a global peer-to-peer network. Smart contracts are self-executing agreements coded directly onto the blockchain, automating processes upon predefined conditions being satisfied. In BCFL, blockchain and smart contracts play a crucial role by providing a secure and transparent framework for decentralized machine learning. Blockchain can ensure the integrity and traceability of data and model updates, while smart contracts can automate the coordination and validation of contributions from different participants. For instance, smart contracts can be used to implement incentive mechanisms for participants in BCFL.
@@ III-C2 Blockchain Tokens
In our BCFL system, a blockchain token named \$FML serves as a quantitative measure of assets. This token can be an indicator of a participant’s trustworthiness within our system. By staking \$FML tokens, participants demonstrate their commitment and reliability. Their tokens will be integrated into the incentive mechanism, as shown in Figure 3. Participants risk losing their staked tokens if they act maliciously. Similar to ERC20 tokens, the \$FML tokens in our BCFL system are fungible and governed by smart contracts. We assume that all participants have acquired sufficient \$FML tokens (e.g., by purchasing \$FML from centralized or decentralized exchanges) before joining the BCFL system.
@@ III-C3 On-Chain Incentive Mechanism
As shown in Figure 3, we follow the existing work [17] and adopt an incentive mechanism to reward honest behaviors and penalize malicious behaviors. As demonstrated in [17], the expected return for malicious participants is negative, and they will ultimately be expelled from the system if they continuously behave maliciously. Different from the existing work [17], which utilizes smart contracts to perform on-chain aggregation of local model updates, our approach allows voters to execute the aggregation off-chain. This improvement can significantly reduce on-chain costs (e.g., gas fees), particularly when dealing with training models that have a large number of parameters.
1. Participants need to stake a fixed amount (i.e., $\delta$) of \$FML tokens into the smart contract to be eligible to participate in the FL training process. 2. In each round, a participant will be randomly selected to act as either an FL training node or a voter by using a verifiable random function (VRF) [19]. 3. A training node is responsible for performing local training using its own data and uploading the local model updates to a public database (e.g., IPFS [^1] https://ipfs.tech/). 4. Voters aggregate all the local model updates for that round to obtain the global model update. They evaluate the global model by running it on their evaluation data to propose voting results, i.e., $\mathsf{vote}_{i}^{r}\in\{\texttt{support},\texttt{oppose}\}$. 5. Based on the majority of the voting result (i.e., $\mathsf{majorVote}^{r}\in\{\texttt{support},\texttt{oppose}\}$), participants in that round will either be rewarded or have their FML tokens slashed. Specifically, if a voter’s choice, $\mathsf{vote}_{i}^{r}$ differs from the majority result (i.e., $\mathsf{vote}_{i}^{t}\not=\mathsf{majorVote}^{r}$), that voter will be penalized; otherwise, they will be rewarded. For all selected training nodes in this round, the majority voting result will determine whether they receive a reward or not. 6. At the beginning of each round, we check the participants’ remaining \$FML token balances in the smart contract: if their balance is still larger than a predefined threshold (i.e., $\delta_{0}$), they will be eligible to continue the following round. Otherwise, they will be removed from the system. 7. For each round $r$, the finalized global model $\theta^{r}$ is determined based on the majority voting result. Specifically: $\theta^{r}=\begin{cases}\hat{\theta}^{r},&\mathsf{majorVote}^{r}=\texttt{support}\\ \theta^{r-1},&\mathsf{majorVote}^{r}=\texttt{oppose}\end{cases}$
@@
Here, $\hat{\theta}^{r}$ represents the aggregated model proposed by voters in round $r$, and $\theta^{r-1}$ is the last confirmed global model.